OpenStudy (anonymous):

domain of f(x)=sqrt(3-x)

OpenStudy (anonymous):

(-infinity,3) <-- bracket or parenthesis?

OpenStudy (anonymous):

parenthesis excludes the 3 because it never actually crosses the x axis right?

OpenStudy (hitaro9):

(-infinity,3] As 3 is included in there (and if its included it does indeed fall in the domain. sqrt(3-3) = sqrt (0) = 0.

OpenStudy (***[isuru]***):

@Hitaro9 is correct!!!

OpenStudy (anonymous):

so interval notation will have a bracket for the domain of sqrt(3-x) = (-infin, 3]

OpenStudy (anonymous):

thanks team

OpenStudy (hitaro9):

No problem. remember that sqrt of 0 is a real number, 0, so anything that gives you that is considered part of the domain.